OpenStudy (anonymous):

Evaluate Lim x˄3-3x˄2+3x-1/x˄3-x x͢ 1

OpenStudy (anonymous):

it is (x-1)^3/x^3-x So, reduced to (x-1)^2/x. substituting 1, the answer is zero.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{x^{3}-3x^2+3x-1}{x^3-x}=\lim_{x \rightarrow 1} \frac{(x-1)^3}{x(x^2-1)}\] \[=\lim_{x \rightarrow 1} \frac{(x-1)^3}{x(x+1)(x-1)}=\lim_{x \rightarrow 1} \frac{(x-1)^2}{x(x+1)}\] Applying limit we find: \[=\frac{(1-1)^2}{1(1+1)}=\frac{0}{2}=0\] Thus the limit of the given function is 0